OpenStudy (anonymous):

The first four terms of a sequence are shown below: 7, 4, 1, -2 Which of the following functions best defines this sequence? f(1) = 7, f(n + 1) = f(n) + 3; for n ≥ 1 f(1) = 7, f(n + 1) = f(n) - 3; for n ≥ 1 f(1) = 7, f(n + 1) = f(n) - 4; for n ≥ 1 f(1) = 7, f(n + 1) = f(n) + 4; for n ≥ 1 I think A?

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

@imqwerty

OpenStudy (anonymous):

Hi anyone know?

OpenStudy (anonymous):

I'm sorry the four terms are 7, 4, 1, -2

OpenStudy (dayakar):

F(n+1)=F(n)+3 if n=1 f(1+1)= F(1) +3 F(2) =7+3 [ since f(1) =7 given] F(2) =10

OpenStudy (anonymous):

so it is A, B, C, or D? Am I right

OpenStudy (dayakar):

no

OpenStudy (dayakar):

B

OpenStudy (anonymous):

ok so it would be subtracting the three?

OpenStudy (dayakar):

yes

OpenStudy (anonymous):

ok thanks

OpenStudy (dayakar):

welcome

OpenStudy (anonymous):

Rushwr do you agree?

OpenStudy (dayakar):

yes,it is

OpenStudy (anonymous):

I'm trying to see if anyone else agrees with the answer before I put it

OpenStudy (anonymous):

ok then thanks again